A manager of Paris Manufacturing, which produces computer hard drives, is planning to lease a new, automated inspection system. The manager believes the new system will be more accurate than the current manual inspection process. The firm has had problems with hard drive defects in the past; the automated system should help catch these defects before the drives are shipped to the final assembly manufacturer. The relevant information follows. Current Manual Inspection System Annual fixed cost = $34,000 Inspection variable cost per unit = $10.00 per unit New, Automated Inspection System Annual fixed cost = $173,000 Inspection variable cost per unit = $0.25 per unit Suppose annual demand is 5,000 units. Should the firm lease the new inspection system? This problem includes break-even calculations. F Round your answer to the nearest whole number. Breakeven volume: hard drives Since the annual demand of 5,000 hard drives is (lower, higher) than the breakeven volume, the firm (should, should not) lease the new inspection system.
Answer #1

Annual Demand D = 5000 units

Cost of Manual Inspection = Fixed Cost + Variable Cost = 34000 + 5000*10 = 84000

Cost of Manual Inspection = Fixed Cost + Variable Cost = 173000 + 5000*0.25 = 174250

Since Cost is much higher for Automated Inspection system, hence firm should not lease for annual demand 5000.

Now we will find out the cost of indifference -

Suppose At X unit, cost will be indifferent.

34000 + X*10 = 173000 + X*0.25

9.75X = 139000

X = 139000/9.75

X = 14256.41

X = 14257 units

Since cost of indifference in is with 14257 units which is higher than 5000 units, hence firm should not lease the new inspection system.